OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

3x+3 =-2x+3+6 3x+3=-2x+9 -3 -3 3x=-2x+6 +2 +2 5x=6 x=\[\frac{ 5 }{ 6 }\]

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

I did it wrong the second 3 should be a 2

OpenStudy (anonymous):

so x=1

OpenStudy (anonymous):

First distribute: 3(x + 1) = −2(x − 1) + 6 3x + 3 = -2x +2 + 6 Now combine like terms 3x + 3 = -2x + (2 + 6) 3x + 3 = -2x + 8 Now isolate the x term +2x +2x ------------------ 5x +3 = 8 -3 -3 ------------- 5x = 5 ------- 5 x=1

OpenStudy (anonymous):

Thank you so much ! :D

OpenStudy (anonymous):

Create your own two-step equation with NO SOLUTION as the solution. Show the steps to solving the equation and explain in words why your work proves that there are NO SOLUTIONS.